OpenStudy (anonymous):

Find the period,range, and amplitude of the cosine function. y= -5 cos 6 x

OpenStudy (xapproachesinfinity):

in general circular function such as \[y=A\cos(Bx+C)+D\] has amplitude A period \[\frac{2\pi}{|B|}\] phase shift c units to left D upward shift

OpenStudy (xapproachesinfinity):

for the range we know that \[-1\le \cos x\le 1\]

OpenStudy (anonymous):

The amplitude is -5 correct?

OpenStudy (anonymous):

Wait or is it just 5?

OpenStudy (xapproachesinfinity):

so we can say that \[-1\le \cos 6x \le 1 \Longrightarrow -5\le -5 \cos 6x \le 5\] \[-5 \le y \le 5\]

OpenStudy (xapproachesinfinity):

oh actually 5 sorry amplitude is absolute value

OpenStudy (anonymous):

Then 1/6pi is the period right?

OpenStudy (anonymous):

I meant 1/3.

OpenStudy (xapproachesinfinity):

2pi/6=pi/3

OpenStudy (xapproachesinfinity):

yea 1/3 pi is good

OpenStudy (anonymous):

Thanks so much for your help!

OpenStudy (xapproachesinfinity):

np :)